# Two-level Nonlinear Preconditioning Methods for Flood Models Posed on Perforated Domains

### v1

@@ 6 Numerical Results
We now proceed with numerical experiments for the proposed algorithms for multiple test cases. These numerical tests will proceed in order of increasing difficulty; they include a porous media equation on an L-shaped domain, a porous media equation on a smaller realistic urban domain, and the Diffusive Wave model on a large realistic urban domain. As mentioned, the fine-scale triangulation is conforming to the coarse polygonal partitioning. For the numerical tests on urban areas (Examples 2 and 3), we ensure that the fine-scale triangulation is identical as the number of subdomains $N$ changes by generating a finer background mesh to be used for various $N$; for example, we can generate a fine-scale triangulation for $N=16\times 16$ subdomains, then use this background triangulation for $N=2\times 2,4\times 4,$ or $8\times 8$ subdomains. Figure 6 shows a matching background triangulation for $N=2\times 4$ and $8\times 16$ subdomains, respectively.
@@
6.1 Porous Media Equation on L-shaped Domain
As a first numerical example, referred to as Example 1, we use an L-shaped domain with a reentering corner. The domain is defined by $D=(-1,1)^{2}$, $\Omega_{S}=(0,1)^{2}$ and $\Omega=D\setminus\overline{\Omega_{S}}$ such that the model domain has singularity/corner in the upper right quadrant of the domain. With this model domain, we consider a porous media equation given by
@@ 6.1 Porous Media Equation on L-shaped Domain
where $\Gamma_{N}=\{(x,y)\in\partial\Omega\,|\,x\neq 1\text{ and }y\neq 1\}$ denotes the Neumann boundary. Consider a piecewise affine finite element discretization and let $\mathbf{u}$ denote a vector of interior node values of size $N_{\Omega}$. We denote the nodal “hat” basis functions $(\eta^{\ell})_{\ell=1}^{N_{\Omega}}$. We define the lumped finite-element mass matrix $M$ and finite element stiffness matrix $A$ such that
$M_{ii}=\sum_{\ell=1}^{N_{\Omega}}\int_{\Omega}\eta_{i}\eta_{\ell}\,\,{\rm d}\mathbf{x},\qquad\text{and}\qquad A_{i\ell}=\int_{\Omega}\nabla\eta_{i}\cdot\eta_{\ell}\,\,{\rm d}\mathbf{x}.$
@@
$F(\mathbf{u})=M\mathbf{u}+A\boldsymbol{\phi},$ (29)
where
$\boldsymbol{\phi}=\max(\mathbf{u},0)^{4}.$
@@
Method Outer Its GMRES Its (per outer it) $N_{it,loc}$
Newton 88 38.1 N/A
Anderson 31 N/A 4.35
2-step 22 35.0 3.96
1-level RASPEN 15 85.3 3.7
2-level RASPEN 6 19.2(+5.3 coarse) 3.94
Table 1: Example 1, $N=9$, initial guess $\mathbf{u}^{0}=0.05$.
Method Outer Its GMRES Its (per outer it) $N_{it,loc}$
Newton 81 33.4 N/A
Anderson 32 N/A 4.64
2-step 16 34.0 4.86
1-level RASPEN 11 53.9 5.08
2-level RASPEN 6 19.2(+4.5 coarse) 4.48
Table 2: Example 1, N=5, initial guess $\mathbf{u}^{0}=0.05$.
Method Outer Its GMRES Its (per outer it) $N_{it,loc}$
Newton 79 33.7 N/A
Anderson 25 N/A 4.36
2-step 13 38.9 6.06
1-level RASPEN 9 35.4 6.96
2-level RASPEN 6 20.5(+3.8 coarse) 5.5
Table 3: Example 1, $N$=3, initial guess $\mathbf{u}^{0}=0.05$.
@@
Convergence curves for this example are shown in Figure 3. Additionally results for this example are summarized in Tables 1, 2, and 3. These tables report the total number of outer iterations and, whenever appropriate, also report the number of GMRES and NRAS iterations. From Figure 3, we see that Newton’s method results in a very large plateau before the region of quadratic convergence is reached. This difference is stark when compared to the other methods. Additionally, for all number of subdomains $N$, two-level RASPEN outperforms the other methods in terms of iteration count, with one-level RASPEN, two-step, and Anderson acceleration following in order. Additionally, we notice that while Anderson does not plateau like Newton’s method, we do not obtain the eventual steep slope of convergence which the other methods achieve.
@@
We see from Tables 1, 2, and 3 that the two-level RASPEN method is consistently resulting in fewer iterations than the other methods regardless of the number of subdomains $N$. The two-level RASPEN method is almost perfectly scalable based on the number of outer and GMRES iterations. We observe that the one-level RASPEN method loses scalability in terms of GMRES iterations as the number of subdomains increases, more so than the other methods which all involve a coarse component of some type. We also observe that Newton’s method and the Two-step method, which both solve a linearized system with the linear two-level RAS preconditioner (with the Trefftz space as the coarse space), result in fair scalability in terms of GMRES iterations. However, both methods still result overall in more GMRES iterations per outer iteration. We additionally remark that the two-step method and Newton’s method result in around the same number of GMRES iterations per outer iteration, with the two-step method resulting in fewer outer iterations than Newton’s method.
6.2 Porous Media Equation on Large Urban Domain
@@
As a second example, referred to as Example 2, we provide numerical results for a Porous Media equation on a large urban domain of size $160\times 160$ metres such that $D=(-80,80)^{2}$ and $\Omega_{S}$ is given by the union of realistic perforations. The buildings are removed from the computational domain, leading to 72 total perforations. The geometry of the buildings have been provided by Métropole Nice Côte d’Azur. After discretization, this model domain contains 35220 triangles and 21317 mesh points. As mentioned, the overlap is set to $\frac{1}{20}\mathcal{H}_{j}$.
With this, we consider a porous media equation given by
@@
where $\Gamma_{N}=\{(x,y)\in\partial\Omega\,|\,x\neq-80\text{ and }y\neq 80\}$ denotes the Neumann boundary; with this, the boundary of each perforation ($\partial\Omega\cap\Omega_{S}$) is included in $\Gamma_{N}$. Here, we take $m=3$ and $c=15$ in (30). The finite element solution for this example is shown in Figure 4.
@@
For this example, we provide numerical results for Newton’s method from Algorithm 1, the two-step method from Algorithm 4, the two-level RASPEN method from Algorithm 3, and the Anderson acceleration method from Algorithm 5. As we have established that one-level RASPEN is not a scalable method for large $N$, we do not provide results for one-level RASPEN in this example. As in the previous example, for the two-step and Newton methods, at each outer iteration, the linear system is solved using GMRES method combined with the linear two-level RAS preconditioner (14).
@@
The results for this example are summarized in Tables 4, 5, 6, and 7 for initial guesses of $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$ and for subdomain partitioning of $N=16\times 16,8\times 8,4\times 4,$ and $2\times 2$, respectively. As mentioned, while the coarse partitionings vary, the background fine-scale triangulation is consistent throughout the numerical experiment. The total number of coarse nodes are given by $N_{\cal V}=$784, 332, 149, and 56 for $N=16\times 16,8\times 8,4\times 4,$ and $2\times 2$, respectively.
@@
Corresponding convergence curves for both initial guesses are provided in Figure 5. Even more than the previous example, Newton’s method has a large plateau in convergence before reaching the region of quadratic convergence. In terms of initial guess, we observe that Newton’s method is much more sensitive to the initial guess than the other methods in terms of outer iteration count. In fact, we see that this change in initial guess can result in a 50% increase in iteration count for Newton’s method. Other than the outer iterations for Newton’s method, there is not much dependence on the initial guess. For all number of subdomains $N$, two-level RASPEN outperforms the other methods in terms of iteration count, with the Two-step, Anderson, and Newton’s methods following in increasing order. We particularly see that the outer iterations of the two-level RASPEN method are extremely robust with respect to $N$, more so than the other accelerated methods.
@@
From Tables 4, 5, 6, and 3, we see again that all two-level algorithms are fairly scalable in terms of GMRES iterations, with the two-level RASPEN method having the fewest GMRES iterations per iteration and the fewest outer iterations. The two-level RASPEN method is almost perfectly scalable in terms of GMRES and otuer iterations. Again, the Two-step method and Newton’s method result in around the same number of GMRES iterations per outer iteration, with the two-step method resulting in fewer outer iterations than Newton’s method. Overall, our conclusions from Example 2 are similar to that of Example 1, with Newton’s method particularly struggling in terms of outer iterations and having the most dependency on the initial guess. However, in terms of solving the linearized system, it appears that the Trefftz coarse space is successful as a component of the linear two-level RAS preconditioner for this nonlinear problem.
@@
Method Its GMRES Its (per it) $N_{it,loc}$
$\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$
Newton 227 346 45.0 46.0 N/A N/A
Anderson 37 52 N/A N/A 4.1 5.07
2-step 25 30 41.1 41.6 4.14 2.83
2-RASPEN 9 8 22.3(+7.4) 19.2 (+9.6) 3.13 2.11
Table 4: Example 2, $N=16\times 16$ subdomains, results for two different initial guesses $\mathbf{u}^{0}=0$ and $\mathbf{u}^{0}=1$.
Method Its GMRES Its (per it) $N_{it,loc}$
$\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$
Newton 227 346 35.2 34.0 N/A N/A
Anderson 35 35 N/A N/A 5.03 5.06
2-step 20 20 38.4 38.1 5.42 4.5
2-RASPEN 7 6 27.1(8.0) 19.5 (8.5) 5.25 3.85
Table 5: Example 2, $N=8\times 8$ subdomains, results for two different initial guesses $\mathbf{u}^{0}=0$ and $\mathbf{u}^{0}=1$.
Method Its GMRES Its (per it) $N_{it,loc}$
$\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$
Newton 227 346 26.2 25.4 N/A N/A
Anderson 22 22 N/A N/A 7.66 7.96
2-step 11 11 30.1 31.9 9.19 9.66
2-RASPEN 7 7 20.4(5.9) 15.3 (6.3) 7.87 6.07
Table 6: Example 2, $N=4\times 4$ subdomains, results for two different initial guesses $\mathbf{u}^{0}=0$ and $\mathbf{u}^{0}=1$.
Method Its GMRES Its (per it) $N_{it,loc}$
$\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$ $\mathbf{u}^{0}=1$ $\mathbf{u}^{0}=0$
Newton 227 346 21.0 18.3 N/A N/A
Anderson 17 17 N/A N/A 10.0 13.93
2-step 8 8 25.6 25.6 16.19 24.53
2-RASPEN 6 6 15.3(4.3) 13.2 (4.3) 13.33 26.21
Table 7: Example 2, $N=2\times 2$ subdomains, results for two different initial guesses $\mathbf{u}^{0}=0$ and $\mathbf{u}^{0}=1$.

### v2

@@ 6 Numerical Results
We now provide numerical experiments for the proposed algorithms (Algorithms 1-5) for multiple test cases. These numerical tests will proceed in order of increasing difficulty; they include a stationary porous medium equation on an L-shaped domain, the similar equation on a small realistic urban domain, and the Diffusive Wave model on a large realistic urban domain.
@@
6.1 Porous Medium Equation on L-shaped Domain
As a first numerical example, referred to as Example 1, we use an L-shaped domain with a reentering corner. The domain is defined by $D=(-1,1)^{2}$, $\Omega_{S}=(0,1)^{2}$ and $\Omega=D\setminus\overline{\Omega_{S}}$ such that the model domain has a singularity/corner in the upper right quadrant of the domain. With this model domain, we consider a porous medium equation given by
@@ 6.1 Porous Medium Equation on L-shaped Domain
where $\Gamma_{N}=\{(x,y)\in\partial\Omega\,|\,x\neq 1\text{ and }y\neq 1\}$ denotes the Neumann boundary. Consider a piecewise affine finite element discretization, and let $\mathbf{u}$ denote a vector of interior node values of size $N_{\Omega}$. We denote the nodal “hat” basis functions $(\eta^{\ell})_{\ell=1}^{N_{\Omega}}$. We define the lumped finite-element mass matrix $M$ and finite element stiffness matrix $A$ such that
$M_{ii}=\sum_{\ell=1}^{N_{\Omega}}\int_{\Omega}\eta_{i}\eta_{\ell}\,\,{\rm d}\mathbf{x},\qquad\text{and}\qquad A_{i\ell}=\int_{\Omega}\nabla\eta_{i}\cdot\nabla\eta_{\ell}\,\,{\rm d}\mathbf{x}.$
@@
$F(\mathbf{u})=M\mathbf{u}+A\max(\mathbf{u},0)^{4}.$ (30)
@@
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 88 38.1 3353 N/A N/A
Anderson 31 N/A N/A 4.35 134.85
2-step 22 35.0 770 3.96 87.12
1-level RASPEN 15 85.3 1280 3.7 55.5
2-level RASPEN 6 19.2(+5.3) 115(+32) 3.94 23.64
Table 1: Example 1, $N=9$, initial guess $\mathbf{u}^{0}=0.05$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 81 33.4 2705 N/A N/A
Anderson 32 N/A N/A 4.64 149.0
2-step 16 34.0 544 4.86 77.8
1-level RASPEN 11 53.9 593 5.08 55.9
2-level RASPEN 6 19.2(+4.5) 115(+27) 4.48 26.9
Table 2: Example 1, $N=5$, initial guess $\mathbf{u}^{0}=0.05$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 79 33.7 2662 N/A N/A
Anderson 25 N/A N/A 4.36 109.0
2-step 13 38.9 506 6.06 78.8
1-level RASPEN 9 35.4 319 6.96 62.6
2-level RASPEN 6 20.5(+3.8) 123(+23) 5.5 33.0
Table 3: Example 1, $N=3$, initial guess $\mathbf{u}^{0}=0.05$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
@@
Convergence curves for this example are shown in Figure 3. Additionally, results for this example are summarized in Tables 1, 2, and 3. These tables report the total number of outer iterations $N_{k}$ and, whenever appropriate, also report the average number of GMRES iterations $N_{\text{GM}}$, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$, the average number of local linear solves on a given subdomain per outer iteration $N_{it,loc}$, and the cumulative number of local linear solves on a given subdomain $N_{k}\cdot N_{it,loc}$.
@@
In terms of outer iterations, from Figure 3, we see that Newton’s method results in a very large plateau before the region of quadratic convergence is reached. This difference is stark when compared to the other methods. Additionally, for all number of subdomains $N$, Two-level RASPEN outperforms the other methods in terms of outer iteration count, with the One-level RASPEN, Two-step, Anderson, and Newton methods following in order. Additionally, we notice that while Anderson does not initially plateau like Newton’s method, we do not obtain the eventual steep slope of convergence which the other methods achieve.
@@
We see from Tables 1, 2, and 3 that the Two-level RASPEN method is scalable in terms of both outer iterations and GMRES iterations, in the sense that these values remain approximately the same for all $N$. In terms of GMRES iterations, we observe that the One-level RASPEN method loses scalability as the number of subdomains increases, more so than the other methods which all involve a coarse component of some type; particularly, as $N$ is doubled in one direction, the GMRES iterations increase and $N_{k}\cdot N_{\text{GM}}$ about doubles as well. We also observe that Newton’s method and the Two-step method, which both solve a linearized system with the linear two-level RAS preconditioner (with the Trefftz space as the coarse space), result in similar values of $N_{\text{GM}}$ for all $N$. However, as the number of Two-step outer iterations is lower than the number of Newton outer iterations, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$ is significantly lower for the Two-step method. The Two-level RASPEN method results in both the smallest value of $N_{k}$ and the smallest value of $N_{\text{GM}}$; therefore, it clearly results in the lowest number of cumulative GMRES iterations $N_{k}\cdot N_{\text{GM}}$. While the Two-level RASPEN has the additional $N_{k}\cdot N_{c}$ coarse linear solves, these coarse solves are generally less expensive than the global GMRES iterations, and the number of coarse linear solves is not exceedingly large.
In terms of local linear solves per subdomain, we do see a slight dependence of $N_{it,loc}$ on the number of subdomains $N$ for all methods, likely due to the fact that the size of the local problems $\frac{N_{\Omega}}{N}$ gets larger as $N$ is decreased. However, between methods for a given $N$, the values of $N_{it,loc}$ appear consistent as expected. Depending on the method, as $N_{k}$ is increased, clearly $N_{k}\cdot N_{it,loc}$ will increase accordingly.
6.2 Porous Medium Equation on Large Urban Domain
@@
As a second example, referred to as Example 2, we provide numerical results for a Porous Medium equation on a large urban domain of size $160\times 160$ meters such that $D=(-80,80)^{2}$ and $\Omega_{S}$ is given by the union of realistic perforations. The buildings are removed from the computational domain, leading to 72 total perforations. The geometry of the buildings have been provided by Métropole Nice Côte d’Azur. After discretization, this model domain contains 35220 triangles and 21317 mesh points. As mentioned, the overlap is set to $\frac{1}{20}\mathcal{H}_{j}$.
For this example, we ensure that the fine-scale triangulation is identical as the number of subdomains $N$ changes by generating a finer background mesh to be used for various $N$; specifically, we can a fine-scale triangulation conforming to the coarse partitioning $N=16\times 16$ subdomains, then use this background triangulation for $N=2\times 2,4\times 4,$ and $8\times 8$ subdomains.
With this, we consider a porous medium equation given by
@@
where $\Gamma_{N}=\{(x,y)\in\partial\Omega\,|\,x\neq-80\text{ and }y\neq 80\}$ denotes the Neumann boundary; with this, the boundary of each perforation ($\partial\Omega\cap\partial\Omega_{S}$) is included in $\Gamma_{N}$. Here, we take $m=3$ and $c=15$ in (31). With a similar finite element discretization to Example 1, the matrix-vector notation of (31) becomes
$F({\bf u})=M{\bf u}+c\,A\max({\bf u},0)^{m},$
with the finite element solution for this example shown in Figure 4.
@@
For this example, we provide numerical results for Newton’s method from Algorithm 1, the Two-step method from Algorithm 4, the Two-level RASPEN method from Algorithm 3, and the Anderson acceleration method from Algorithm 5. As we have established that One-level RASPEN is not a robust method for large $N$, we do not provide results for One-level RASPEN in this example. As in the previous example, for the Two-step and Newton methods, at each outer iteration, the linear system is solved using GMRES method combined with the linear two-level RAS preconditioner (14).
@@
The results for this example are summarized in Tables 4, 5, 6, and 7 for initial guesses of $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$ and for subdomain partitioning of $N=16\times 16,8\times 8,4\times 4,$ and $2\times 2$, respectively. While the coarse partitionings vary, the background fine-scale triangulation is consistent throughout the numerical experiment. The total number of coarse nodes are given by $N_{\cal V}=$784, 332, 149, and 56 for $N=16\times 16,8\times 8,4\times 4,$ and $2\times 2$, respectively. These tables report the total number of outer iterations $N_{k}$ and, whenever appropriate, also report the average number of GMRES iterations $N_{\text{GM}}$, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$, the average number of local linear solves on a given subdomain per outer iteration $N_{it,loc}$, and the cumulative number of local linear solves on a given subdomain $N_{k}\cdot N_{it,loc}$.
@@
Corresponding convergence curves for both initial guesses are provided in Figure 5. In terms of outer iterations, even more than the previous example, Newton’s method has a large plateau in convergence before reaching the region of quadratic convergence. For all number of subdomains $N$, Two-level RASPEN outperforms the other methods in terms of outer iteration count, with the Two-step, Anderson, and Newton’s methods following in increasing order. We particularly see that the outer iterations of the Two-level RASPEN method are extremely robust with respect to $N$, more so than the other accelerated methods. Additionally, from Tables 4, 5, 6, and 7, we see that there is a slight increase in $N_{k}$ for the Two-step and Anderson methods as $N$ is increased, but the increase in $N_{k}$ is not proportional to the increase in $N$. The number of outer iterations $N_{k}$ stays extremely consistent for the Newton and Two-level RASPEN methods as $N$ is varied. In terms of initial guess, we observe that Newton’s method is much more sensitive to the initial guess than the other methods in terms of outer iteration count. In fact, we see that this change in initial guess can result in a 50% increase in iteration count for Newton’s method. Other than the outer iterations for Newton’s method, there is no significant dependence on the initial guess.
From Tables 4, 5, 6, and 7, we see that $N_{\text{GM}}$ for the Two-level RASPEN method is significantly smaller than that of the Two-step and Newton methods. Additionally, as the number of outer iterations $N_{k}$ is the lowest for the Two-level RASPEN method, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$ is the lowest for the Two-level RASPEN method. As in the previous example, the Two-step method and Newton’s method result in around the same number of GMRES iterations per outer iteration $N_{\text{GM}}$, with this value slightly increasing as $N$ is increased; however, this increase is not proportional to the increase in $N$. As the Two-step method results in fewer outer iterations $N_{k}$ than Newton’s method, the total number of GMRES iterations $N_{k}\cdot N_{\text{GM}}$ is significantly lower for the Two-step method. This can result in between 10 and 20 times more total GMRES iterations for Newton’s method.
In terms of local linear solves per subdomain, our conclusions are similar to those of Example 1. Between methods, the values of $N_{it,loc}$ appear fairly consistent, but depending on $N_{k}$, $N_{k}\cdot N_{it,loc}$ will increase accordingly.
@@
Overall, our conclusions from Example 2 are similar to that of Example 1, with Newton’s method particularly struggling in terms of outer iterations and having an additional dependency on the initial guess. However, in terms of solving the linearized system, it appears that the Trefftz coarse space is successful as a component of the linear two-level RAS preconditioner for this nonlinear problem.
@@
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 45.0 10215 N/A N/A
Anderson 37 N/A N/A 4.1 152
2-step 25 41.1 1028 4.14 104
2-RASPEN 9 22.3(+7.4) 201(+67) 3.13 28
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 46.0 15916 N/A N/A
Anderson 52 N/A N/A 5.07 264
2-step 30 41.6 1248 2.83 85
2-RASPEN 8 19.2(+9.6) 154(+77) 2.11 17
Table 4: Example 2, $N=16\times 16$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 35.2 7990 N/A N/A
Anderson 35 N/A N/A 5.03 176
2-step 20 38.4 768 5.42 108
2-RASPEN 7 27.1(+8.0) 190(+56) 5.25 37
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 34.0 11764 N/A N/A
Anderson 35 N/A N/A 5.06 177
2-step 20 38.1 762 4.5 90
2-RASPEN 6 19.5(+8.5) 117(+51) 3.85 23
Table 5: Example 2, $N=8\times 8$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 26.2 5947 N/A N/A
Anderson 22 N/A N/A 7.66 169
2-step 11 30.1 331 9.19 101
2-RASPEN 7 20.4(+5.9) 143(+41) 7.87 55
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 25.4 8788 N/A N/A
Anderson 22 N/A N/A 7.96 175
2-step 11 31.9 351 9.66 106
2-RASPEN 7 15.3(+6.3) 107(+44) 6.07 43
Table 6: Example 2, $N=4\times 4$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.
$\mathbf{u}^{0}=1$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 227 21.0 4767 N/A N/A
Anderson 17 N/A N/A 10.0 170
2-step 8 25.6 205 16.19 130
2-RASPEN 6 15.3(+4.3) 92(+26) 13.33 80
$\mathbf{u}^{0}=0$
Method $N_{k}$ $N_{\text{GM}}$ $N_{k}\cdot N_{\text{GM}}$ $N_{it,loc}$ $N_{k}\cdot N_{it,loc}$
Newton 346 18.3 6332 N/A N/A
Anderson 17 N/A N/A 13.93 237
2-step 8 25.6 205 24.53 196
2-RASPEN 6 13.2(+4.3) 79(+26) 26.21 157
Table 7: Example 2, $N=2\times 2$ subdomains, results for initial guesses $\mathbf{u}^{0}=1$ and $\mathbf{u}^{0}=0$. Values in brackets represent $N_{c}$ for the $N_{GM}$ column and $N_{k}\cdot N_{c}$ for the $N_{k}\cdot N_{GM}$ column.